OpenStudy (anonymous):

For the following problem, y varies directly as the square of x. If y = 8 when x = 2, find y when x is 1.

hartnn (hartnn):

so, \(\Large y \alpha x^2\) let the constant of proportionality be 'k' so, \(\Large y= k x^2\) when y=8, x= 2 can you find 'k' first ?

OpenStudy (anonymous):

k=4

OpenStudy (anonymous):

no 2

hartnn (hartnn):

are you sure ? \(8 = k\times 2^2 \\ 8 = 4k \\\)

hartnn (hartnn):

yes! k =2 is correct :)

OpenStudy (anonymous):

I forgot about the exponent

hartnn (hartnn):

now we have \(\large y = 2x^2\) just plug in x=1 !

OpenStudy (anonymous):

y=2?

hartnn (hartnn):

thats correct :)

OpenStudy (anonymous):

Thank you! :)

hartnn (hartnn):

welcome ^_^